a total of 242 mosquitoes filled the room, but only females bite. If the ratio of males to females is four to five, how many of the mosquitoes could bite?
Answer by addingup(3677)

Ratio: 4:5, means that out of every 4+5=9 mosquitoes, 4 are male and 5 female.
242/9 = 26.89 There are 26.89 groups of 9 mosquitoes
26.89*4 = 107.56 male and
242-107.56 = 134.44 female
-------------------------
I don't know how they told you to present the answer. Since you end up with fractions and any time you talk about a living organism (including us)you have to make allowances for the fact that you can't cut 0.56 of a human being.